The radius of the earth is about 3959 miles.
The distance from the Reading traffic circle to the Natick Center is about 1/3 mile.
Find the angle in seconds made by two rays from the center of the earth through these two landmarks.
(1 radian = 180/π = 57.29577951 degrees, and 1 second = 1/3600th of one degree)

Answer by stanbon(75887)

The radius of the earth is about 3959 miles.
The distance from the Reading traffic circle to the Natick Center is about 1/3 mile.
Find the angle in seconds made by two rays from the center of the earth through these two landmarks.
(1 radian = 180/π = 57.29577951 degrees, and 1 second = 1/3600th of one degree)
-------------------
arc length = theta*radius
----
(1/3)mile = theta*3959
---
theta = (1/3)/3959 = 0.000084196.. radians
-----------------------------------------------
Convert to degrees:
0.000084196(180/pi) = 0.004824 degrees
= (0.004824)(60 seconds)
= 0.28945 seconds
==========================
Cheers,
Stan H.